Assume that there is no storage or other carrying costs/value for corn. Assume that corn is not perishable, that is it can be stored and used later and that the continuous compounded risk-free rate is 20% per year.

If a 6-month corn future currently costs $4.75 a bushel and the spot price currently costs $4.50 explain what portfolio a trader should construct to take advantage of these prices.
Answer #1

first we will calculate theotrical futures price

F=Se^rt

where s is spot price

r is risk free rate

= 4.5*e^0.2*6/12 =4.9732

but actual futures price is 4.75 so futures are underpriced

arbitrage portfolio is long futures short in cash market and lend money